OpenStudy (anonymous):

HELP PLEASE !!! Which point lies on the line with point-slope equation y - 2 = 5(x - 6)?

OpenStudy (anonymous):

is it (-6, 2)??

OpenStudy (johnweldon1993):

not quite...to check....Try plugging it in...plug in (-6) for x and (2) for y

OpenStudy (anonymous):

(6, -2)

OpenStudy (johnweldon1993):

seeing as how -2 - 2 is -4....and 6 - 6 is 0....no...that is not an option either

OpenStudy (johnweldon1993):

You see what I am doing correct? I'm simply plugging in the numbers from the choice you give...and seeing if the equation works

OpenStudy (anonymous):

is (6, 2) an option?

OpenStudy (anonymous):

soooooooo(-6, -2)

OpenStudy (johnweldon1993):

Lets plug it in 2 - 2 = 5(6 - 6)? 0 = 5(0) 0 = 0 yes it is....(6,2)

OpenStudy (anonymous):

thank you(: sorry Im kind of slow..lol

OpenStudy (johnweldon1993):

don't worry about it....that's what this site is about is helping people understand :)